In triangleABC, AC = 3x-14, BC= 7x-74, and measure of angle C= 90 degrees. Find the value of x that makes triangleABC a right isosceles triangle.
Answer by checkley75(3666)

3X-14=7X-74
3X-7X=-74+14
-4X=-60
X=-60/-4
X=15 ANS.
PROOF
3*15-14=7*15-74
45-14=105-74
31=31